An employee who produces x units per hour earns an hourly wage of y=0.55 + 7 (in dollars). Find the hourly wage for an employee who produces 13 units per hour. Please help me understand this. Thanks so much!
Answer by checkley71(8403)

Y=0.55+7
IF YOU MEANT Y=0.55(X)+7 X BEING THE NUMBER OF UNITS PRODUCED. THEN:
Y=.55*13+7
Y=7.15=7
Y=14.15 DOLLARS PER HOUR FOR PRODUCING 13 UNITS PER HOUR IS THE ANSWER.
IF YOU MEANT Y=0.55+7(X). THEN
Y=0.55+7*13
Y=0.55+91
Y=91.55 DOLLARS PER HOUR. [I'LL TAKE THIS JOB!!!!!]